# SMRU: Split-and-Merge Recurrent-based UNet for Acoustic Echo Cancellation and Noise Suppression

### v1

@@ Abstract
Index Terms:
Acoustic echo cancellation, noise suppression.
I Introduction
@@
Figure 1: Overview diagram of the proposed hybrid AEC system.
Figure 2: Architecture of the proposed SMRU. Different modules are indicated with different colors for better illustrations. (a) Overall diagram of the proposed SMRU. (b) Detail structure of the multi-scale band split layer. (c) Detail structure of the band merge layer. (d) Detail structure of the variable frame rate block. (e) Detail structure of the inter-band MLP.
@@
The rest of the paper is organized as follows. Sec. II presents the proposed framework. Sec. III and Sec. IV namely give the experimental setups and results. Some conclusions are drawn in Sec. V.
II Proposed Method
@@ II-A Split and Merge
II-A1 Multi-scale band split layer
@@
${\tilde{R}}_{p}=Cat(Conv_{K=k_{p1},S=s_{p}}(R_{p}),\cdots,Conv_{K=k_{pM},S=s_{p}}(R_{p})),\vspace{-0.3cm}$ (2)
${H}=Norm(Conv(Merge(\tilde{R}_{1},\cdots,\tilde{R}_{P}))),\vspace{-0.05cm}$ (3)
@@
II-A2 Band merge layer
@@
$\left\{U_{1},\cdots,U_{Q}\right\}=Sub\text{-}band\text{-}Split\left(U\right),\vspace{-0.1cm}$ (4)
${G}_{q}=MLP_{q}\left(Norm\left(U_{q}\right)\right),\vspace{-0.05cm}$ (5)
${G}=Merge\left({G}_{1},\cdots,{G}_{Q}\right),\vspace{-0.1cm}$ (6)
$\hat{S}^{\left(1\right)}=\sum_{i}^{8}I_{i}\otimes G_{i},\vspace{-0.2cm}$ (7)
@@
II-B Recurrent UNet
II-B1 Variable frame rate block
@@
${Z}_{\downarrow}=TimeDownSample(Z_{i}),\vspace{-0.1cm}$ (8)
${Z}_{\downarrow}^{(1)}=Reshape(FC(GRU(Norm(Z_{\downarrow})))+Z_{\downarrow}),\vspace{-0.1cm}$ (9)
${Z}_{\downarrow}^{(2)}=Inter\text{-}band\text{-}MLP({Z}_{\downarrow}^{(1)})+{Z}_{\downarrow}^{(1)},\vspace{-0.1cm}$ (10)
$Z_{o}=TimeUpSample({Z}_{\downarrow}^{(2)}),\vspace{-0.1cm}$ (11)
@@
II-B2 Causal time Down-Sample and Up-Sample layer
@@
II-B3 Inter-band MLP shuffler
@@
II-B4 Cross-scale skip connections
In addition to the standard skip connections in UNet, we also introduce cross-scale skip connections, as shown by the blue curve in Figure 2. We adopt a connection mode similar to dense connections [20], but instead of channel-wise concatenation, we sum them after normalization. By doing so, nearly no extra computational overhead is introduced. We observe that the strategy can effectively improve performance, which will be revealed in Sec. IV-B.
II-C Post-processing module
@@ II-D Loss function
$\mathcal{L}_{vad}=10\log_{10}(||\hat{S}\times(1-{\mathbb{I}}_{vad})||_{2}^{2}+\epsilon),\vspace{-0.2cm}$ (13)
@@
$\mathcal{L}=\mathcal{L}_{MAE}+0.1\mathcal{L}_{echo}+\beta\mathcal{L}_{vad},\vspace{-0.2cm}$ (14)
where $\beta$ is set to balance the capability between echo suppression and near-end speech preservation, whose impact will be shown in Sec. IV-B.
III Experimental setup
III-A Data preparation
@@ III-C Evaluation metrics
TABLE I: The objective results of proposed SMRU and different baselines in the test set. BOLD indicates the best score.
@@
Figure 3: AECMOS metrics of the blind test set under the DT scenario.
Figure 4: Spectrum visualizations of an example. (a) Mix audio. (b) Target near-end speech. (c) Estimated spectrum processed by DeepFilterNet. (d) Estimated spectrum processed by SMRU-S.
IV Experimental results
IV-A Result comparisons with baselines
We compare the proposed SMRU with five advanced baselines on the test set, and quantitative results are shown in Table I. Four modes of SMRU are investigated, namely tiny (T), small (S), large (L), and huge (H), with the complexity varying from 50 M/s to 6.83 G/s, in terms of MACs. Baseline methods include NSNet [25], DTLN [8], DeepFilterNet, FastFullSubNet, and BSRNN. The latter three models were originally proposed for the speech enhancement task, and we adapt them to AEC task by using the same input as the proposed method. The real-time factor (RTF) is measured on an Intel Core (TM) i7-9750H CPU clocked at 2.60 GHz. From the table, several observations can be made. First, with the increase in computational complexity, the objective metric scores of the proposed method are gradually improved, where the tiny version achieves overall better performance over NSNet but only with around one-third percentage in complexity. Moreover, although DTLN performs well in the ST-FE scenario, it lacks capability in other scenarios. In DT and ST-NE scenarios, DTLN performs worse compared to SMRU-S, which has only a quarter of its complexity. For the large version, SMRU further outperforms BSRNN and FastFullSubNet with less complexity. It fully validates the superiority of the proposed method. Besides, when a PostNet is adopted, notable improvements can be observed in both DT and ST-NE cases and just slight degradation in ERLE for the ST-FE case, which validates the effectiveness of the post-processing. Finally, compared with BSRNN and FastFullSubNet, SMRU-L enjoys a notably lower RTF, which can be attributed to the proposed UNet structure with a multi-level time sampling strategy.
@@ IV-A Result comparisons with baselines
TABLE II: Ablation study on the proposed SMRU-S. Cond. 1 represents the condition of using a multi-scale band split layer, while Cond. 2 represents the condition of using cross-scale skip connections.
@@
TABLE III: Ablation study on the weight $\beta$ of the proposed VAD-oriented loss.
@@
IV-B Ablation study
Ablation studies are conducted on SMRU-S to investigate the effects of using multi-scale band split layer and cross-scale skip connections in the test set. For the case without the multi-scale band split layer, the single-scale convolution is used for band splitting, i.e., the number of convolutions in the convolution set for each region is set to 1. The results in Table II show that removing the multi-scale band split layer can lead to significant performance degradation, as the single-scale convolution cannot provide frequency representations at different resolutions. Besides, the cross-scale skip connections can also provide a notable performance improvement by only introducing minimal computational overhead.
Table III shows the performance of different $\beta$ values in the loss function. When $\beta=0.0002$, both PESQ in DT and ST-NE scenarios and ERLE in ST-FE scenario show an improvement over $\beta=0$. However, when $\beta$ further increases, the model exhibits better capability in far-end echo suppression, i.e., a higher ERLE score, but at the cost of more speech distortion, i.e., lower PESQ and SI-SNR scores. Due to space constraints, we do not traverse more $\beta$ options, and $\beta=0.0002$ seems adequate to well balance between echo cancellation and target speech preservation.
V Conclusion

### v2

@@ Abstract
Index Terms—  Acoustic echo cancellation, noise suppression
1 Introduction
@@
Fig. 1: Overview diagram of the proposed hybrid AEC system.
Fig. 2: Architecture of the proposed SMRU. Different modules are indicated with different colors for better illustrations. (a) Overall diagram of the proposed SMRU. (b) Detail structure of the multi-scale band split layer. (c) Detail structure of the band merge layer. (d) Detail structure of the variable frame rate block. (e) Detail structure of the inter-band MLP.
@@
The rest of the paper is organized as follows. Sec. 2 presents the proposed framework. Sec. 3 and Sec. 4 namely give the experimental setups and results. Some conclusions are drawn in Sec. 5.
2 Proposed Method
@@ 2.1 Split and Merge
2.1.1 Multi-scale band split layer
@@
${\tilde{R}}_{p}=Cat(Conv_{K=k_{p1},S=s_{p}}(R_{p}),\cdots,Conv_{K=k_{pM},S=s_{p}}(R_{p})),\vspace{-0.0cm}$ (2)
${H}=Norm(Conv(Merge(\tilde{R}_{1},\cdots,\tilde{R}_{P}))),\vspace{-0.00cm}$ (3)
@@
2.1.2 Band merge layer
@@
$\left\{U_{1},\cdots,U_{Q}\right\}=Sub\text{-}band\text{-}Split\left(U\right),\vspace{-0.0cm}$ (4)
${G}_{q}=MLP_{q}\left(Norm\left(U_{q}\right)\right),\vspace{-0.0cm}$ (5)
${G}=Merge\left({G}_{1},\cdots,{G}_{Q}\right),\vspace{-0.0cm}$ (6)
$\hat{S}^{\left(1\right)}=\sum_{i}^{8}I_{i}\otimes G_{i},\vspace{-0.0cm}$ (7)
@@
2.2 Recurrent UNet
2.2.1 Variable frame rate block
@@
${Z}_{\downarrow}=TimeDownSample(Z_{i}),\vspace{-0.0cm}$ (8)
${Z}_{\downarrow}^{(1)}=Reshape(FC(GRU(Norm(Z_{\downarrow})))+Z_{\downarrow}),\vspace{-0.0cm}$ (9)
${Z}_{\downarrow}^{(2)}=Inter\text{-}band\text{-}MLP({Z}_{\downarrow}^{(1)})+{Z}_{\downarrow}^{(1)},\vspace{-0.0cm}$ (10)
$Z_{o}=TimeUpSample({Z}_{\downarrow}^{(2)}),\vspace{-0.0cm}$ (11)
@@
2.2.2 Causal time Down-Sample and Up-Sample layer
@@
2.2.3 Inter-band MLP shuffler
@@
2.2.4 Cross-scale skip connections
In addition to the standard skip connections in UNet, we also introduce cross-scale skip connections, as shown by the blue curve in Figure 2. We adopt a connection mode similar to dense connections [20], but instead of channel-wise concatenation, we sum them after normalization. By doing so, nearly no extra computational overhead is introduced. We observe that the strategy can effectively improve performance, which will be revealed in Sec. 4.2.
2.3 Post-processing module
@@ 2.4 Loss function
$\mathcal{L}_{vad}=10\log_{10}(||\hat{S}\times(1-{\mathbb{I}}_{vad})||_{2}^{2}+\epsilon),\vspace{-0.0cm}$ (13)
@@
$\mathcal{L}=\mathcal{L}_{MAE}+0.1\mathcal{L}_{echo}+\beta\mathcal{L}_{vad},\vspace{-0.0cm}$ (14)
where $\beta$ is set to balance the capability between echo suppression and near-end speech preservation, whose impact will be shown in Sec. 4.2.
3 Experimental setup
3.1 Data preparation
@@ 3.3 Evaluation metrics
Table 1: The objective results of proposed SMRU and different baselines in the test set. BOLD indicates the best score.
@@
Fig. 3: AECMOS metrics of the blind test set under the DT scenario.
Fig. 4: Spectrum visualizations of an example. (a) Mix audio. (b) Target near-end speech. (c) Estimated spectrum processed by DeepFilterNet. (d) Estimated spectrum processed by SMRU-S.
4 Experimental results
4.1 Result comparisons with baselines
We compare the proposed SMRU with five advanced baselines on the test set, and quantitative results are shown in Table 1. Four modes of SMRU are investigated, namely tiny (T), small (S), large (L), and huge (H), with the complexity varying from 50 M/s to 6.83 G/s, in terms of MACs. Baseline methods include NSNet [25], DTLN [8], DeepFilterNet, FastFullSubNet, and BSRNN. The latter three models were originally proposed for the speech enhancement task, and we adapt them to AEC task by using the same input as the proposed method. The real-time factor (RTF) is measured on an Intel Core (TM) i7-9750H CPU clocked at 2.60 GHz. From the table, several observations can be made. First, with the increase in computational complexity, the objective metric scores of the proposed method are gradually improved, where the tiny version achieves overall better performance over NSNet but only with around one-third percentage in complexity. Moreover, although DTLN performs well in the ST-FE scenario, it lacks capability in other scenarios. In DT and ST-NE scenarios, DTLN performs worse compared to SMRU-S, which has only a quarter of its complexity. For the large version, SMRU further outperforms BSRNN and FastFullSubNet with less complexity. It fully validates the superiority of the proposed method. Besides, when a PostNet is adopted, notable improvements can be observed in both DT and ST-NE cases and just slight degradation in ERLE for the ST-FE case, which validates the effectiveness of the post-processing. Finally, compared with BSRNN and FastFullSubNet, SMRU-L enjoys a notably lower RTF, which can be attributed to the proposed UNet structure with a multi-level time sampling strategy.
@@ 4.1 Result comparisons with baselines
4.2 Ablation study
Ablation studies are conducted on SMRU-S to investigate the effects of using multi-scale band split layer and cross-scale skip connections in the test set. For the case without the multi-scale band split layer, the single-scale convolution is used for band splitting, i.e., the number of convolutions in the convolution set for each region is set to 1. The results in Table 2 show that removing the multi-scale band split layer can lead to significant performance degradation, as the single-scale convolution cannot provide frequency representations at different resolutions. Besides, the cross-scale skip connections can also provide a notable performance improvement by only introducing minimal computational overhead.
Table 3 shows the performance of different $\beta$ values in the loss function. When $\beta=0.0002$, both PESQ in DT and ST-NE scenarios and ERLE in ST-FE scenario show an improvement over $\beta=0$. However, when $\beta$ further increases, the model exhibits better capability in far-end echo suppression, i.e., a higher ERLE score, but at the cost of more speech distortion, i.e., lower PESQ and SI-SNR scores. Due to space constraints, we do not traverse more $\beta$ options, and $\beta=0.0002$ seems adequate to well balance between echo cancellation and target speech preservation.
Table 2: Ablation study on the proposed SMRU-S. Cond. 1 represents the condition of using a multi-scale band split layer, while Cond. 2 represents the condition of using cross-scale skip connections.
@@ 4.2 Ablation study
Table 3: Ablation study on the weight $\beta$ of the proposed VAD-oriented loss.
@@
5 Conclusion